OpenStudy (anonymous):

When we're talking about things like sin(x) What does that 'x' represent? Does it represent a radian value?

OpenStudy (anonymous):

Sin, tan, cos..etc are trig functions. Essentially the 'x' represents the angle. So it can be in either radians or degrees.

OpenStudy (anonymous):

Oh, and if the radians value and degree value are equivalent to each other, like: \[\pi/3 = 30 degrees\], plugging either of this in a trig function such as sin result in 1/2

OpenStudy (anonymous):

Like @Zelda said, x represents the angle (be it degrees or radians), while sinx, cosx, tanx, etc., are ratios of the length of one side to the length of another. (The "sides" I'm referring to belong to the triangle containing the angle x.)

OpenStudy (anonymous):

But if we have the value as a greek letter, we know that it is in radians, correct? such as \[\sin \alpha \cos \beta \tan \theta\]

OpenStudy (anonymous):

Nope, not necessarily.

OpenStudy (anonymous):

Theta is just generally used to represent an angle. Not necessarily in radians or degrees. The other two letters are rarely, if ever used.

OpenStudy (anonymous):

@zelda so when we solve it and find \[\theta\], how would we know if we have the value in radians or degrees? Like, solving it through taking the arcsin, for example.

OpenStudy (anonymous):

@reviewer, well generally if you use a graphing or even scientific calculator, there is a radian or degree mode. Depending on what mode you'e in, you'll get the answer in that form.

OpenStudy (anonymous):

By convention, greek letters indicate radians, but that's not always the case.

OpenStudy (anonymous):

*usually* indicate radians